# Matteo Escudé

### v5

@@ 2 Model
There is an unknown binary state of the world $\omega$, taking values in $\{0,1\}$. A sender and a receiver initially assign prior probability $\overline{p}\in[0,1]$ to $\omega=1$. The sender can acquire information about the state by choosing a signal and observing its realization. Information acquisition is covert in the sense that the receiver observes neither the sender’s chosen signal, nor its realization. Formally, a feasible signal is a pair $(\mathcal{S},\pi)$ where $\mathcal{S}$ is a finite set of possible realizations and $\pi\in\Delta(\{0,1\}\times\mathcal{S})$ is a joint distribution such that the marginal distribution of the state coincides with the prior $\overline{p}$. [^16] We will soon restrict $\mathcal{S}$ to be a subset of $[0,1]$. See Remark 1.
After acquiring information, the sender communicates with the receiver. Her acquired private information determines what she is able to say; formally, if the sender’s chosen signal is $(\mathcal{S},\pi)$ and its realization is $s\in\mathcal{S}$, she must send a message from the finite set $M(\mathcal{S},\pi;s)$ to the receiver. I refer to the mapping $M$, assigning sets $\{M(\mathcal{S},\pi;s)\}_{s\in\mathcal{S}}$ to every signal $(\mathcal{S},\pi)$, as the verifiability structure, which is common knowledge among the two players. $M$ describes how the sender’s acquired private information constrains how she can communicate with the receiver. In doing so, it also captures what the sender can prove about what she learnt to the receiver. [^17] As discussed in the introduction, I do not take a stance on where a given verifiability structure comes from, but rather characterize outcomes in the large class of structures described in Definition 1. The constraints captured by $M$ may arise for several reasons. For example, it may be that the receiver (or a third-party) can check the sender’s claims, or that the sender has reputational concerns, or that what she can say is dictated by disclosure regulation.
@@
The preferences of the receiver are not modelled explicitly as they play no role in the analysis. The receiver can be interpreted either as a literal player, who takes a payoff-relevant action following the sender’s message, or as a ‘market’ in which the sender obtains return $v(p)$ at belief $p$ [^18] In Section 5 I characterize receiver-optimal verifiability under the (mild) assumption that the receiver values information about the state.
@@ 2.1 Beliefs determine messages
I restrict attention to verifiability structures in which beliefs determine messages, defined as follows.
@@
$M$ is such that beliefs determine messages if for any pair of signals $(\mathcal{S},\pi)$ and $(\mathcal{S}^{\prime},\pi^{\prime})$, if $\pi(\cdot|s)=\pi^{\prime}(\cdot|s^{\prime})$ then $M(\pi,\mathcal{S};s)=M(\pi^{\prime},\mathcal{S}^{\prime};s^{\prime})$ for any $s\in\mathcal{S}$, $s^{\prime}\in\mathcal{S}^{\prime}$.
If beliefs determine messages it is only the information about the state contained in the signal realization—that is: the sender’s own private posterior belief—that determines what messages the sender can use.
The underlying assumption is that the sender cannot directly prove anything about how she chose to acquire information (that is: about the signal $(\pi,\mathcal{S})$), but only about what she learned regarding the state. Observe that proving information about the state indirectly proves something about the sender’s choice of signal. That is, if a certain message $m$ proves to the receiver that the sender’s belief lies in some set $S_{m}\subseteq[0,1]$, the sender is also proving that she acquired information in a way that assigns positive probability to her obtaining a belief in $S_{m}$. [^19] Note, however, that this information is payoff-irrelevant, as the sender’s payoff only depends on the receiver’s posterior about the state. The beliefs determine messages assumption imposes that such a message $m$ proves nothing more than this to the receiver.
@@
In contrast, information about the state can more easily be verified, for example by a trusted third party (e.g. by an auditor, a fact-checker or an accreditation body) who certifies what the sender reports about the payoff-relevant state. Importantly, such certification typically does not coincide with full disclosure of what the sender discovers: it may involve coarse grades or thresholds, and perhaps leave the sender the possibility of not making a report in case of bad news. Section 2.2 below provides more detailed examples of such verifiability structures.
Importantly, abstracting away from this secondary layer of communication has a methodological payoff, as it allows to study the problem using a belief-based approach.
@@
When restricting attention to verifiability structures in which beliefs determine messages, it is without further loss to restrict the sender to choosing signals such that $\mathcal{S}\subseteq[0,1]$ and $\pi(1|s)=s$ for all $s\in\mathcal{S}$. With this normalization, the signal realization $s$ coincides with the sender’s own private posterior belief about the state. I shall sometimes refer to $s$ as the sender’s (interim) type. When beliefs determine messages, only the sender’s private posterior belief determines what she can communicate, so we can lighten notation further by omitting the dependence of $M$ on the signal. Henceforth $M$ will simply denote a mapping assigning to each $s\in[0,1]$ a finite set of messages $M(s)$.
2.1.1 Technical assumptions
The following technical assumptions on the verifiability structure $M$ are maintained. Let $\mathcal{M}\equiv\cup_{s\in[0,1]}M(s)$ denote the set of all possible messages. $M^{-1}(m)\subseteq[0,1]$ is assumed to be closed for all $m\in\mathcal{M}$. The map $s\mapsto\max_{m\in M(s)}\min M^{-1}(m)$, $s\in[0,1]$ is assumed to be upper semi-continuous. [^20] A sufficient condition for this continuity requirement is that the set of all possible messages $\mathcal{M}$ is finite. These technical assumptions guarantee equilibrium existence, established in Appendix A.
@@ 2.2 Examples of verifiability structures
Suppose the sender is forced to reveal whatever she has discovered about the state. This could be due to a deliberate (ex ante) choice by the sender (e.g. by hiring a trusted third-party to certify that no acquired information is withheld) or to stringent disclosure regulation. Formally, we say $M$ is the mandatory disclosure verifiability structure if $M(s)=\{s\}$ for each $s\in[0,1]$, capturing that the sender is forced to truthfully report her belief about the state. It is immediate (for more details, see the discussion in Section 4), that this verifiability structure allows the sender to attain Kamenica and Gentzkow (2011)’s ‘full-commitment’ value in equilibrium. $\diamond$
@@
Mandatory disclosure as described in the previous example is a special case of a verifiability structure in which the sender is able to prove precisely what she discovered to the receiver. More generally, the sender might have the option of doing so, but may also be able to misreport or conceal her findings. For example, she might hire a trusted third-party certifier, but retain the option of preventing him from making any disclosure. [^21] We can describe this particular situation with the verifiability structure $M(s)=\{s,m_{0}\}$ for all $s\in[0,1]$. Upon learning $s$, the sender can either allow the third-party certifier to make a (truthful) disclosure of $s$ or can prevent him from reporting anything (sending the null message $m_{0}$).
@@
The following environments exhibit instead partial verifiability: $M(s)\subseteq M(s^{\prime})$ for some $s,s^{\prime}\in[0,1]$, meaning that type $s^{\prime}$ can always imitate type $s$.
@@
At the other end of the spectrum from full verifiability, we have cheap talk situations, where the sender can claim anything regardless of what she discovered. This could be a consequence of the information acquisition technology: for example, it could be that learning only produces information which is ‘soft’ or easy-to-falsify (e.g. when using eyewitness accounts or anonymous sources). Formally, we have that $M(s)=\bar{M}$ for each $s\in[0,1]$, for some given—perhaps very large—finite set $\bar{M}$. Section 3.3 discusses outcomes in environments of this kind. $\diamond$
@@
To formalize this notion, let $\mathcal{P}$ denote a partition of $[0,1]$; a partitional verifiability structure $M_{\mathcal{P}}$ assigns to each $P\in\mathcal{P}$ a message $m_{P}$. That is: if type $s$ lies in the partition element $P\subseteq[0,1]$, she reveals the partition element using message $m_{P}$. In applications, partitional structures are typically of the interval form: each $P\in\mathcal{P}$ is an interval. In this case, the verifiability structure can be interpreted as a scoring protocol, with a higher score corresponding to a higher (non-overlapping) interval of types. Remark 2 provides a general result for comparing such structures from the perspective of the sender. $\diamond$
@@
The sender might be able to make certifiable claims of the form ‘I discovered the state is high at least with probability $s$’. This report is credible provided no type $s^{\prime}<s$ can also make such a claim. A special case of this is when the sender privately learns via a binary test and she can prove she obtained good news (e.g. she can prove her posterior is above the prior), but can always claim to have obtained no information. [^22] For example, there could be two messages only: the ‘null’ message $m_{0}\in M(s)$ for all $s\in[0,1]$ and the ‘good news’ message $m_{G}\in M(s)$ for all $s\geq\bar{p}$.
Verifiability structures of this kind can be described by certifiable thresholds. Formally, they are defined by a sequence of sender types, $0=s_{0}<s_{1}<\dots<s_{N}\leq 1$, and a set of messages $\{m_{0},\dots,m_{N}\}$ such that $m_{i}\in M(s)$ iff $s\geq s_{i}$. In this environment, the sender can prove to the receiver that a threshold $s_{i}$ (corresponding to a probability $s_{i}$ of $\omega=1$) has been reached. She can always claim to have attained a lower threshold, but not a higher one. One such example is solved in detail in Section 3.1 and a (simple, 2-message) structure in this class proves to be receiver-optimal (Section 5). $\diamond$
@@ 2.3 Strategies and solution concept
A strategy for the sender is a pair $(\pi,\mu)$. $\pi$ is the signal chosen by the sender; $\mu$ is her messaging mixed strategy: $(\pi,s)\mapsto\mu(\pi,s)$ assigns to every $(\pi,s)$ an element of $\Delta(M(s))$. Denote by $\beta:\mathcal{M}\to[0,1]$ the receiver’s belief about the state.
The solution concept is in the spirit of perfect Bayesian equilibrium: the sender acquires and transmits information optimally given how the receiver forms beliefs and the receiver’s posterior belief about the state must be consistent with the verifiability structure on and off the equilibrium path.
More formally, given the triple $(\pi,\mu,\beta)$, say that type $s\in[0,1]$ is on-path iff $s\in\operatorname*{supp}\pi_{S}$, where $\pi_{S}$ denotes the marginal distribution of the signal. Message $m\in\mathcal{M}$ is on-path iff $\mu(m|\pi,s)>0$ for some $s\in[0,1]$ on-path. Consistency with the verifiability structure (on and off the equilibrium path) requires that $\beta(m)\in\operatorname*{conv}M^{-1}(m)$ for all $m\in\mathcal{M}$, where $\operatorname*{conv}M^{-1}(m)$ denotes the convex hull of $M^{-1}(m)$. For $m\in\mathcal{M}$ on-path, $\beta(m)$ is further required to satisfy Bayes’ rule. Appendix A contains a formal definition of the equilibrium concept and an existence proof.
3 Equilibrium skepticism and unraveling
Recall that—as discussed in the previous section—we restrict attention to non-decreasing and upper semi-continuous payoff functions $v$ and verifiability structures $M$ in which beliefs determine messages (Definition 1) and which satisfy the technical assumptions in Section 2. We will refer to parameters in this class as admissible.
@@ 3 Equilibrium skepticism and unraveling
Sender type $s$ is lowest-consistent with some message $m$ if no lower type can send $m$ [^23] In the terminology of Seidmann and Winter (1997), sender type $s$ is lowest-consistent with some message $m$ if it is the lowest of the worst-case types for $M^{-1}(m)$. Type $s$ is therefore able to separate from every lower type, no matter what the receiver believes, by sending message $m$.

### v6

@@ 2 Model
There is an unknown binary state of the world $\omega$, taking values in $\{0,1\}$. A sender and a receiver initially assign prior probability $\overline{p}\in[0,1]$ to $\omega=1$. The sender can acquire information about the state by choosing a signal and observing its realization. Information acquisition is covert in the sense that the receiver observes neither the sender’s chosen signal, nor its realization. Formally, a feasible signal is a pair $(\mathcal{S},\pi)$ where $\mathcal{S}$ is a finite set of possible realizations and $\pi\in\Delta(\{0,1\}\times\mathcal{S})$ is a joint distribution such that the marginal distribution of the state coincides with the prior $\overline{p}$: $\sum_{s\in\mathcal{S}}\pi(1,s)=\overline{p}$. [^13] We will soon restrict $\mathcal{S}$ to be a subset of $[0,1]$, see Remark 1.
After acquiring information, the sender communicates with the receiver. Her acquired private information determines what she is able to say: if the sender’s chosen signal is $(\mathcal{S},\pi)$ and its realization is $s\in\mathcal{S}$, she must send a message from the non-empty finite set $M(\mathcal{S},\pi;s)$ to the receiver. I refer to the mapping $M$, assigning sets $(M(\mathcal{S},\pi;s))_{s\in\mathcal{S}}$ to every signal $(\mathcal{S},\pi)$, as the verifiability structure, which is common knowledge among the two players. $M$ describes how the sender’s acquired private information constrains how she can communicate with the receiver. In doing so, it also captures what the sender can prove about what she learned to the receiver.
@@
The preferences of the receiver are not modeled explicitly as they play no role in the analysis. The receiver can be interpreted either as a literal player, who takes a payoff-relevant action following the sender’s message, or as a ‘market’ in which the sender obtains return $v(p)$ at belief $p$.
@@ 2.1 Beliefs determine messages
Operationalizing the notion of verifiability structure requires articulating how the sender’s private information shapes what she can say and, vice versa, how the receiver’s beliefs are restricted following a given message from the sender. I do so by restricting attention to verifiability structures in which beliefs determine messages, defined as follows.
@@
$M$ is such that beliefs determine messages iff for every pair of signals $(\mathcal{S},\pi)$ and $(\mathcal{S}^{\prime},\pi^{\prime})$, and every $s\in\mathcal{S}$, $s^{\prime}\in\mathcal{S}^{\prime}$, if $\pi(\cdot|s)=\pi^{\prime}(\cdot|s^{\prime})$ then $M(\mathcal{S},\pi;s)=M(\mathcal{S}^{\prime},\pi^{\prime};s^{\prime})$.
If beliefs determine messages, it is only the information about the state contained in the signal realization—that is: the sender’s own private posterior belief—that determines what messages the sender can use.
The underlying assumption is that the sender cannot directly prove anything about how she chose to acquire information (that is: about the signal $(\mathcal{S},\pi)$), but only about what she learned regarding the state. Observe that proving information about the state indirectly proves something about the sender’s choice of signal. That is, if a certain message $m$ proves to the receiver that the sender’s belief lies in some set $S_{m}\subseteq[0,1]$, the sender is also proving that she acquired information in a way that assigns positive probability to her obtaining a belief in $S_{m}$. The beliefs determine messages assumption imposes that such a message $m$ proves nothing more than this to the receiver.
@@
In contrast, information about the state can more easily be verified, for example by a trusted third party (such as an auditor, a fact-checker or an accreditation body) who certifies what the sender reports about the payoff-relevant state. Importantly, such certification typically does not coincide with full disclosure of what the sender discovers: it may involve coarse grades or thresholds, and perhaps leave the sender the possibility of redacting a report if it reveals bad news. Section 2.2 below provides more detailed examples of such verifiability structures.
Importantly, abstracting away from this secondary layer of communication has a methodological payoff, as it allows the problem to be studied using a belief-based approach.
@@
Given the restriction to verifiability structures in which beliefs determine messages, it follows that:
  1. (i)
    It is without further loss to restrict the sender to choosing signals taking values in $[0,1]$ and such that each realization coincides with the sender’s own private posterior belief about the state (i.e., such that $\pi(1|s)=s$ when $s\in[0,1]$ realizes). I shall frequently refer to $s$ as the sender’s type.
  2. (ii)
    We can lighten notation further by omitting the dependence of $M$ on the signal $(\mathcal{S},\pi)$. Henceforth $M$ will simply denote a mapping assigning to each $s\in[0,1]$ a non-empty finite set of messages $M(s)$.
Before turning to examples, I state a closure assumption on the verifiability structures considered, which is maintained throughout the analysis. For a given verifiability structure $M$, let $\mathcal{M}\equiv\cup_{s\in[0,1]}M(s)$ denote the set of all possible messages. For any given message $m\in\mathcal{M}$, let
$M^{-1}(m)\equiv\{s\in[0,1]:m\in M(s)\}$
denote the set of sender types that can send message $m$. We assume that this set always contains its infimum:
$\inf M^{-1}(m)\in M^{-1}(m)\quad\text{for every}\quad m\in\mathcal{M}.$ (1)
Therefore, for any given message, there is a well-defined lowest type that could have sent it. This condition ensures that the receiver’s ‘maximally skeptical’ belief—which will be used throughout the analysis and is based on this lowest type—is always well-defined.
@@ 2.2 Examples of verifiability structures
Suppose the sender is forced to reveal whatever she has discovered about the state. This could be due to a deliberate (ex ante) choice by the sender (e.g., by hiring a trusted third-party to certify that no acquired information is withheld) or to stringent disclosure regulation. Formally, we say $M$ is the mandatory disclosure verifiability structure if $M(s)=\{s\}$ for each $s\in[0,1]$, capturing that the sender is forced to truthfully report her belief about the state. It is immediate (for more details, see the discussion in Section 4) that this verifiability structure allows the sender to attain Kamenica and Gentzkow (2011)’s ‘full-commitment’ value in equilibrium.  $\diamond$
@@
Mandatory disclosure as described in the previous example is a special case of a verifiability structure in which the sender is able to prove precisely what she discovered to the receiver. More generally, the sender might have the option of doing so, but may also be able to misreport or conceal her findings. For example, she might hire a trusted third-party certifier, but retain the option of preventing him from making any disclosure. [^14] We can describe this particular situation with the verifiability structure $M(s)=\{s,m_{\emptyset}\}$ for each $s\in[0,1]$. Therefore, upon learning $s$, the sender can either allow the third-party certifier to make a (truthful) disclosure of $s$ or can prevent him from reporting anything, sending the null message $m_{\emptyset}$.
@@
The following environments exhibit instead partial verifiability in the sense that $M(s)\subseteq M(s^{\prime})$ for some $s,s^{\prime}\in[0,1]$, meaning that type $s^{\prime}$ can always imitate type $s$.
@@
At the other end of the spectrum from full verifiability, we have cheap talk situations, where the sender can claim anything regardless of what she discovered. This could be a consequence of the information acquisition technology: for example, it could be that learning only produces information which is ‘soft’ or easy-to-falsify. Formally, we have that $M(s)=\bar{M}$ for each $s\in[0,1]$, for some given—perhaps very large—finite set $\bar{M}$. Section 3.2 discusses outcomes in environments of this kind.  $\diamond$
@@
To formalize this notion, let $\mathcal{P}$ denote a partition of $[0,1]$; a partitional verifiability structure $M_{\mathcal{P}}$ assigns to each $P\in\mathcal{P}$ a message $m_{P}$. That is: if type $s$ lies in the partition element $P\subseteq[0,1]$, she must reveal the partition element she belongs to by sending message $m_{P}$. In applications, partitional structures are typically of the interval form: each $P\in\mathcal{P}$ is an interval. In this case, the verifiability structure can be interpreted as a scoring protocol, with a higher score corresponding to a higher (non-overlapping) interval of types. Remark 2 provides a general result for comparing such structures from the perspective of the sender.  $\diamond$
Example 5 (Dye verifiability).
Suppose that the sender can always make a verifiable claim about what she discovers, unless she learns nothing, in which case she has no evidence to present. Suppose also that, regardless of what she learns, she always has the option of suppressing her verifiable information by claiming that she learned nothing instead. This simple verifiability structure was introduced by Dye (1985) and Jung and Kwon (1988) to show that unraveling can fail when verifiability is partial. In my model, this situation can be formalized by setting $M(s)=\{s,m_{\emptyset}\}$ for each $s\neq\overline{p}$, while $M(\overline{p})=\{m_{\emptyset}\}$. Message $m_{\emptyset}$ can therefore be interpreted as the ‘non-disclosure’ message. [^15] To illustrate how unraveling fails with this verifiability structure in the case of a fixed (exogenous) signal, consider a strictly increasing $v$ and a signal such that with probability $\alpha\in(0,1)$ the sender learns the value of the state and with complementary probability $1-\alpha$ learns nothing. This signal gives rise to three possible sender types: 0, $\overline{p}$ and $1$. In equilibrium, sender type 0 will pool with type $\overline{p}$ by sending message $m_{\emptyset}$, while type $1$ will separate by sending message $1$. This signal cannot, in fact, be an equilibrium in my model, as it is subject to a deviation which worsens the meaning of $m_{\emptyset}$. See Theorem 2 and the discussion in the continuation of Example 5 on 1 for more details.
This is a special case of a more general class of verifiability structures in which there is a set $S_{v}\subseteq[0,1]$ of verifiable sender types such that $M(s)=\{s,m_{\emptyset}\}$ for $s\in S_{v}$ while $M(s)=\{m_{\emptyset}\}$ for $s\notin S_{v}$.  $\diamond$
@@
The sender might be able to make certifiable claims of the form ‘I discovered the state is high at least with probability $s$’. This report is credible provided no type $s^{\prime}<s$ can also make such a claim. A special case of this is when the sender privately learns via a binary test and she can prove she obtained good news (e.g., she can prove her posterior is above the prior), but can always claim to have obtained no information.
Verifiability structures of this kind can be described by certifiable thresholds. Formally, they are defined by a sequence of sender types, $0=s_{0}<s_{1}<\dots<s_{N}\leq 1$, and a set of messages $\{m_{0},\dots,m_{N}\}$ such that $m_{i}\in M(s)$ iff $s\geq s_{i}$. In this environment, the sender can prove to the receiver that a threshold $s_{i}$ (corresponding to a probability $s_{i}$ of $\omega=1$) has been reached. She can always claim to have attained a lower threshold, but not a higher one. One such example was solved in detail in Section 1.1.  $\diamond$
@@ 2.3 Strategies and solution concept
In light of Remark 1, rather than writing $(\mathcal{S},\pi)$ for the sender’s choice of signal, it creates no ambiguity to suppress the set of realizations and only denote it by the (finite-support) distribution $\pi$, where it is understood that the implied set of signal realizations is
$\mathcal{S}=\operatorname*{supp}\pi_{S}=\{s\in[0,1]:\pi(0,s)+\pi(1,s)>0\},$
where $s\mapsto\pi_{S}(s)=\pi(0,s)+\pi(1,s)$ denotes the marginal distribution of the signal.
We can therefore denote a strategy for the sender by a pair $(\pi,\mu)$, where $\pi$ is the signal chosen and $\mu$ is her messaging mixed strategy: $(\pi,s)\mapsto\mu(\pi,s)$ assigns to every $(\pi,s)$ an element of $\Delta(M(s))$. Denote by $\beta:\mathcal{M}\to[0,1]$ the receiver’s posterior belief about the state. Given the triple $(\pi,\mu,\beta)$, say that type $s\in[0,1]$ is on-path iff $s\in\operatorname*{supp}\pi_{S}$ and that message $m\in\mathcal{M}$ is on-path iff
$m\in\operatorname*{supp}\mu(\cdot|\pi,s)=\{m\in M(s):\mu(m|\pi,s)>0\}$
for some $s\in[0,1]$ on-path.
The solution concept is in the spirit of perfect Bayesian equilibrium: the sender acquires and transmits information optimally given how the receiver forms beliefs and the receiver’s posterior belief about the state is formed using Bayes’ rule on the equilibrium path and is consistent with the verifiability structure off-path. As the sender’s optimality conditions and the receiver’s on-path Bayesian updating are standard, their formal statement is relegated to Appendix A. Off-path, consistency of the receiver’s belief with the verifiability structure requires that
$\beta(m)\in\operatorname*{conv}M^{-1}(m)\quad\text{for every}\quad m\in\mathcal{M},$ (2)
where $\operatorname*{conv}M^{-1}(m)$ denotes the convex hull of $M^{-1}(m)$.
[…]
Finally, the following lemma provides a continuity condition on the verifiability structure ensuring the existence of an equilibrium.
Lemma 1 (Existence).
If $s\mapsto\max_{m\in M(s)}\min M^{-1}(m)$ for $s\in[0,1]$ is upper semi-continuous, an equilibrium exists.
The proof is in Appendix A. The condition guarantees that the sender, when facing a skeptical receiver assigning belief $\min M^{-1}(m)$ to every message $m$, has an optimal choice of signal. A sufficient condition for this continuity requirement is that the set of all possible messages $\mathcal{M}$ is finite.
Since this assumption is not necessary to state the results in the next section, it is not maintained. It will be explicitly invoked in Section 4, where equilibrium existence is required to state the comparative statics results.
3 Unpersuasive and unraveling equilibria
Recall that—as discussed in the previous section—we restrict attention to non-decreasing and upper semi-continuous payoff functions $v$ and verifiability structures $M$ in which beliefs determine messages (Definition 1) and such that (1) holds. We will refer to parameters in this class as admissible.
@@ 3 Unpersuasive and unraveling equilibria
That is, sender type $s\in[0,1]$ is lowest-consistent with message $m$ if no lower type can send $m$. Type $s$ can therefore separate from every lower type—regardless of the receiver’s belief—by sending mm. [^16] In the terminology of Seidmann and Winter (1997), sender type $s$ is lowest-consistent with some message $m$ if it is the lowest of the worst-case types for $M^{-1}(m)$. The maintained assumption that beliefs determine messages further implies that the receiver’s belief about $\omega=1$ following message $m$ must be at least $s$ (both on- and off-path) as captured by equilibrium condition (2).
Note that Definition 2 does not rely on the verifiability structure $M$ being such that beliefs determine messages. [^17] It does, however, rely on the signal realizations being in $[0,1]$ and on the sets of available messages depending only on the signal realization (and not directly on the signal). As discussed in Remark 1, these restrictions entail no loss of generality once $M$ is assumed to be such that beliefs determine messages. If beliefs do not determine messages, however, even if sender type $s\in[0,1]$ is lowest-consistent with message $m$, the receiver could hold—off-path—any belief about the state following $m$. This is because the sender faces no constraints on the joint distribution of state and signal, other than the requirement that the state marginal must equal the prior. So, even if the receiver’s belief about the sender’s type would still have to be supported on types at least equal to $s$, this would not translate into any restriction on her belief about the state.
The following terminology will also prove useful in stating the results.
Definition 3.
Given a sender strategy $(\pi,\mu)$ and receiver belief $\beta$, say that the receiver is maximally skeptical following message $m$ iff $\beta(m)=\min M^{-1}(m)$. Say that the sender reveals all acquired information iff for every $s\in\operatorname*{supp}\pi_{S}$ and every $m\in\operatorname*{supp}\mu(\cdot|\pi,s)$, $\beta(m)=s$.
Equipped with these definitions, we can turn to studying the properties of equilibria.
Definition 4 (Unpersuasive equilibria).
An equilibrium $(\pi,\mu,\beta)$ is unpersuasive iff $v(\beta(m))=v(\overline{p})$ for every message $m$ on-path.
In unpersuasive equilibria, any variation in the receiver’s belief is payoff-irrelevant for the sender: she obtains the same payoff she would achieve without information acquisition and transmission, with probability one. The first result shows that, in equilibria which are not unpersuasive, the sender’s behavior and outcomes are tightly pinned down by the primitives of the model.
Theorem 1 (Unraveling equilibria).
Suppose that $(\pi,\mu,\beta)$ is an equilibrium which is not unpersuasive. Then:
  1. (i)
    Every on-path sender type $s$ is lowest-consistent with every $m\in\operatorname*{supp}\mu(\cdot|\pi,s)$.
  2. (ii)
    The receiver is maximally skeptical following every on-path message.
  3. (iii)
    The sender reveals all acquired information.
We call such equilibria unraveling.
The proof is in Appendix B. Notice that Theorem 1 (i) is saying two things at once. First, it states that only types that are lowest-consistent with some message can be on-path. The verifiability structure $M$ therefore restricts the set of possible outcomes of the sender’s equilibrium information acquisition to those that can then be credibly communicated to the receiver (i.e., to types that are lowest-consistent with some message). Second, it is saying that every on-path sender type will, indeed, only use messages she is lowest-consistent with in equilibrium.
Note also that the theorem is not making the stronger claim that the equilibrium set can be partitioned into unpersuasive and unraveling equilibria, as an equilibrium may be both. [^18] To illustrate, adjust the example in Section 1.1 by increasing the prior to $\overline{p}=0.5$. It can be verified that there is an equilibrium in which the sender acquires no information and sends message $m_{G}$ on-path. This equilibrium is both unpersuasive and unraveling.
The logic behind the theorem is driven by the fact that, if the receiver were not maximally skeptical following an on-path message, the sender could deviate to a signal ‘worsening the meaning’ of that message, i.e., a signal such that the message ends up being sent by a lower type. This is profitable for the sender as the set of on-path messages is unchanged, but probability is shifted towards messages that lead to higher payoffs.
This deviation arises whenever the receiver’s equilibrium beliefs vary in a way that affects the sender’s payoff—i.e., in equilibria that are not unpersuasive. It is precisely this variation in payoff that gives the sender an incentive to shift probability from ‘bad news’ messages to ‘good news’ ones. This temptation evidently vanishes when all on-path messages yield the same payoff.
Notice how this reasoning relies on the assumption that beliefs determine messages, which aligns the receiver’s skepticism about the sender’s type with her skepticism about the state. If type $s$ is lowest-consistent with message $m$, the sender cannot arbitrarily worsen its meaning, as she must believe that $\omega=1$ with probability at least $s$ when sending $m$. Consequently, the receiver’s belief about the state following $m$ cannot fall below $s$, as implied by equilibrium condition (2).
Finally, this intuition also hinges on the monotonicity and binary state assumptions, as they ensure that worsening the meaning of a message never corresponds to a decrease in the probability that a ‘good news’ message is sent. Outside of the binary-state case with monotone payoff, this need not be the case.
3.1 The sender’s value in unraveling equilibria
Theorem 1 also directly implies that the sender obtains the same expected payoff across all unraveling equilibria. This expected payoff can be expressed as the concave envelope of a skepticism-adjusted version of $v$. Letting
$\underline{v}_{M}(s)\equiv v\left(\max_{m\in M(s)}\min M^{-1}(m)\right),$ (3)
for each $s\in[0,1]$, we have the following corollary of Theorem 1.
Corollary 1 (Unique value of persuasion).
In every unraveling equilibrium the sender’s expected payoff is $(\operatorname*{cav}\underline{v}_{M})(\overline{p})$.
That is, the sender’s equilibrium expected payoff is the concave envelope of the ‘value of interim sender beliefs’, when facing a receiver who is maximally skeptical following every message. In the example from Section 1.1, these objects are depicted in Figure 1(b): $\underline{v}_{M}$ is the solid black line, while $\operatorname*{cav}\underline{v}_{M}$ is the thick dotted line and $(\operatorname*{cav}\underline{v}_{M})(\overline{p})$ is the black dot.
Proof of Corollary 1.
Consider any unraveling equilibrium $(\pi,\mu,\beta)$ and define $w_{\beta}(s)\equiv\max_{m\in M(s)}v(\beta(m))$, $s\in[0,1]$. The sender’s equilibrium expected payoff is $(\operatorname*{cav}w_{\beta})(\overline{p})$, since—for a given $\beta$ and optimally chosen messages—the sender faces a canonical Bayesian persuasion problem with value of posteriors $w_{\beta}$. Since $w_{\beta}\geq\underline{v}_{M}$, we have that $(\operatorname*{cav}w_{\beta})(\overline{p})\geq(\operatorname*{cav}\underline{v}_{M})(\overline{p})$.
Observe next that $\underline{v}_{M}=w_{\beta}$ on $\operatorname*{supp}\pi_{S}$. This holds because, for every $s\in\operatorname*{supp}\pi_{S}$, Theorem 1 (i) implies that $s$ is lowest-consistent with some message, so that $\underline{v}_{M}(s)=v(s)$ by construction and Theorem 1 (iii) implies that $w_{\beta}(s)=v(s)$. It follows that, in an auxiliary Bayesian persuasion problem in which the sender faces value of posteriors $\underline{v}_{M}$, the sender can obtain an expected payoff of $(\operatorname*{cav}w_{\beta})(\overline{p})$ by choosing signal $\pi$. Since $(\operatorname*{cav}\underline{v}_{M})(\overline{p})$ is the highest expected payoff the sender can attain in this problem, it must be that $(\operatorname*{cav}w_{\beta})(\overline{p})\leq(\operatorname*{cav}\underline{v}_{M})(\overline{p})$. ∎
3.2 Proving news better than the prior
We have so far established that any equilibrium which is not unpersuasive must be unraveling, meaning that it exhibits the properties described in Theorem 1. We now provide a sufficient condition on the primitives of the model which ensures that all equilibria are unraveling.